Moving Coframes

Peter Olver Department of Mathematics University of Minnesota Minneapolis, MN 55455, USA Email: olver@ima.umn.edu

In this talk I will present a new approach to the theory of moving frames that allows one to explicitly compute them for arbitrary Lie group and pseudogroup actions. Applications to differential invariants, equivalence, and computer vision, as well as computer algebra implementation issues will be discussed.